# Geometry of convex geometries

### v1

@@ 4.3 Convex dimension
Edelman and Jamison [11] provided a nice characterization of convex geometries via order convexity. Given a universe $U$ of size $n$ and a total order $e_{1}\leq e_{2}\leq\ldots\leq e_{n}$, call an ending interval of $\leq$ any set of the form $\{e\in U:e_{i}\leq e\}$ for some $e_{i}\in U$. Given a set $\Upsilon=\{\leq_{1},\ldots,\leq_{d}\}$ of $d$ total orders on $U$, we say that a set family $\mathcal{C}\subseteq 2^{U}$ is generated by the set $\Upsilon$ if $\varnothing\in\mathcal{C}$ and a nonempty set $C$ belongs to $\mathcal{C}$ if and only if $C$ is the intersection of $d$ ending intervals, one from each order $\leq_{i}$ of $\Upsilon$, i.e., if there exists $(e_{i})_{1\leq i\leq d}\in U^{d}$ such that $C=\{e\in U:e_{i}\leq_{i}e,\forall 1\leq i\leq d\}$. Edelman and Jamison proved the following result:
Theorem 18 ([11, Theorem 5.2]).
@@
An ideal $I$ of a poset is a subset that is downwards-closed, i.e. if $x\leq y$ and $y\in I$, then $x\in I$. From Theorem 19 by definition we get a justification of the name ideals of convex geometries:
@@ 4.5 Examples of ideals of convex geometries
Edelman and Jamison [11] and Korte et al. [21] presented numerous examples of convex geometries and antimatroids, arising from geometry, language theory, and chip firing games. Bandelt et al. [2] presented simplicial complexes and median set systems as examples of conditional antimatroids. We will show below that multisimplicial complexes and median set systems are in fact ideals of ideal alignments, a particular class of convex geometries.
A simplicial complex $L$ on a set $U$ is a family of subsets of $U$, called simplices or faces of $L$, such that if $\sigma\in L$ and $\sigma^{\prime}\subseteq\sigma$, then $\sigma^{\prime}\in L$. The facets of $L$ are the maximal (by inclusion) faces of $L$. The dimension $d$ of $L$ is the size of its largest face. A multi-subset $\sigma$ of $U$ is a subset of $U$ such that each element $e\in U$ is given with its multiplicity $n_{\sigma}(e)$ (the number of times, $e$ occurs in $\sigma$). A multisimplicial complex $L$ on a set $U$ is a family of multi-subsets of $U$, such that if $\sigma\in L$ and $\sigma^{\prime}$ is a multi-subset of $U$ such that $n_{\sigma^{\prime}}(e)\leq n_{\sigma}(e)$ for each $e\in U$, then $\sigma^{\prime}\in L$. The size of a face $\sigma$ of $L$ is the size of $\{e\in U:n_{\sigma}(e)>0\}$. The dimension $d$ of $L$ is the size of its largest face.
A median set system is a set family $(U,\mathcal{C})$ satisfying (${}^{\prime})$, (C2), and
@@
The resulting meet-semilattice is called a median semilattice [1, 4, 25]. Median semilattices are locally lower distributive. Indeed, by a result of Birkhoff and Kiss [4] they can be characterized by the property that all intervals $[\hat{0},x]$ are distributive lattices and three elements have an upper bound whenever each pair of them does.
@@
Let $\leq$ be a partial order on $U$. The ideal alignment ${\mathcal{D}}_{P}$ [11] of the poset $P=(U,\leq)$ consists of all ideals of $P$. It was noticed in [11] that the ideal alignments are convex geometries and it was shown in [11, Theorem 3.2] that a convex geometry $\mathcal{A}=(U,\mathcal{C})$ is an ideal alignment if and only if $\mathcal{C}$ is union-closed. From the lattice point of view, ideal alignments simply correspond to distributive lattices.
We say that a pair $\mathcal{A}=(U,\mathcal{C})$ is an ideal semialignment if $\mathcal{C}$ contains $\varnothing$, is intersection-closed, locally extendable in the sense that if $X,Y\in\mathcal{C}$ such that $X\subseteq Y$, then there is $e\in Y\setminus X$ such that $X\cup e\in\mathcal{C}$, and locally union-closed, in the sense that if $X,Y\in\mathcal{C}$ and there exists $Z\in\mathcal{C}$ such that $X,Y\subseteq Z$, then $X\cup Y\in\mathcal{C}$. The next result shows that ideal semialignments are ideals of ideal alignments, hence in particular ideals of convex geometries:
Proposition 22.
Every ideal semialignment $\mathcal{A}=(U,\mathcal{C})$ with $k$ maximal convex sets is an ideal of an ideal alignment $\mathcal{A}^{+}$ such that $\operatorname{VC-dim}(\mathcal{A}^{+})\leq k\cdot\operatorname{VC-dim}(\mathcal{C})$.
@@
We will prove that $\mathcal{A}=(U,\mathcal{C})$ is an ideal of an ideal alignment by induction on the number of inclusion-maximal convex sets in $\mathcal{C}$. If there is only one such set, then $\mathcal{A}$ is a convex geometry and we are done. Otherwise, take a maximal set =$\mathcal{M}=\{X_{1},\ldots,X_{k}\}$ of at least two maxima such that $\bigcap_{i=1}^{k}X_{i}=M$ is maximal. Now define $\mathcal{C}^{+}$ by adding all sets of the form $Y_{1}\cup\ldots\cup Y_{k}$ with $M\subseteq Y_{i}\subseteq X_{i}$ for $1\leq i\leq k$ to $\mathcal{C}$. We set $\mathcal{A}^{+}=(U,\mathcal{C}^{+})$.
First notice that $\mathcal{C}^{+}$ has a new maximal element $\bigcup_{i=1}^{k}X_{i}$, while none of the sets in $\mathcal{M}$ is maximal anymore. Hence, $\mathcal{C}^{+}$ has less inclusion-maximal convex sets than $\mathcal{C}$. Let us show that $\mathcal{C}$ is an ideal of $\mathcal{C}^{+}$. For this purpose let $X$ be an inclusion-maximal convex set from $\mathcal{C}$ and $R=Y_{1}\cup\ldots\cup Y_{k}\subseteq X$ with $R\in\mathcal{C}^{+}$. Then for all $1\leq i\leq k$ we have $Y_{i}\subseteq X_{i}\cap X$. But by the maximality in the choice of $\mathcal{M}$ this implies $X_{i}\cap X=M$ and hence $X\in\mathcal{M}$, say $X=X_{k}$. But then $Y_{i}=M$ for $11\leq i\leq k-$ 1 and hence $R=Y_{k}\in\mathcal{C}$.
@@
Now, we show that $\mathcal{C}^{+}$ is intersection-closed. Let $A,B\in\mathcal{C}^{+}$. If $A,B\in\mathcal{C}$, then there is nothing to show, so suppose first $A\in\mathcal{C}$ and $B\in\mathcal{C}^{+}\setminus\mathcal{C}$. Hence we can represent $B=Y_{1}\cup\ldots\cup Y_{k}$ and we consider $A\cap(Y_{1}\cup\ldots\cup Y_{k})=(A\cap Y_{1})\cup\ldots\cup(A\cap Y_{k})$. Since all of these sets are subsets of $A$, since $\mathcal{C}$ is locally union-closed, their union is in $\mathcal{C}$. If $A,B\in\mathcal{C}^{+}\setminus\mathcal{C}$, then $A=Y_{1}\cup\ldots\cup Y_{k}$ and $B=Y^{\prime}_{1}\cup\ldots\cup Y^{\prime}_{k}$ and $(Y_{1}\cup\ldots\cup Y_{k})\cap(Y^{\prime}_{1}\cup\ldots\cup Y^{\prime}_{k})=Y_{1}\cap(Y^{\prime}_{1}\cup\ldots\cup Y^{\prime}_{k})\cup\ldots Y_{k}\cap(Y^{\prime}_{1}\cup\ldots\cup Y^{\prime}_{k})$ but for each $i$ here we have $M\subseteq Y_{i}\cap(Y^{\prime}_{1}\cup\ldots\cup Y^{\prime}_{k})\subseteq X_{i}$, hence their union is in $\mathcal{C}^{+}$ by definition of $\mathcal{C}^{+}$.
@@
In order to show that $\mathcal{C}^{+}$ is locally union-closed let $A,B\in\mathcal{C}^{+}$ and let $A,B\subseteq Z$. Since we can assume that $A\notin\mathcal{C}$, without loss of generality we can set $Z=\bigcup_{i=1}^{k}X_{i}$ and $A=Y_{1}\cup\ldots\cup Y_{k}$ as before. Now, $B=(B\cap X_{1})\cup\ldots\cup(B\cap X_{k})$, where each term is in $\mathcal{C}$. Then, $A\cup B=(Y_{1}\cup(B\cap X_{1})\cup\ldots\cup Y_{k}\cup(B\cap X_{k}))$, where each term is in $\mathcal{C}$ by local union-closedness. Hence by definition of $\mathcal{C}^{+}$, $A\cup B\in\mathcal{C}^{+}$.
@@
Finally, let us show that $\mathcal{C}^{+}$ is locally extendable. For this purpose let $A,B\in\mathcal{C}^{+}$ distinct sets such that $B$ is a maximal convex set and $A\subseteq B$. If $B\in\mathcal{C}$, then since $\mathcal{C}$ is an ideal, also $A\in\mathcal{C}$ and there is nothing to show. So suppose $A\in\mathcal{C}^{+}$ and $B\in\mathcal{C}^{+}\setminus\mathcal{C}$. We know that $B=\bigcup_{i=1}^{k}X_{i}$ if $A\in\mathcal{C}$, then if $A\cap X_{i}\subsetneq X_{i}$ for some $i$, then there is an element $e\in X_{i}\setminus A\subseteq B\setminus A$ such that $(A\cap X_{i})\cup e\in\mathcal{C}$ since $\mathcal{C}$ is locally extendable. Now, $\mathcal{C}^{+}$ is locally union-closed also $A\cup(A\cap X_{i})\cup e=A\cup e\in\mathcal{C}^{+}$. If $A\in\mathcal{C}^{+}\setminus\mathcal{C}$, then $A=Y_{1}\cup\ldots\cup Y_{k}$ as usual and $Y_{i}\subsetneq X_{i}$ for some $i$. Since $\mathcal{C}$ is locally extendable, there is $e\in X_{i}\setminus Y_{i}\subseteq B\setminus A$ such that $Y_{i}\cup e\in\mathcal{C}$. But then by definition of $\mathcal{C}^{+}$, we have that $A\cup e\in\mathcal{C}^{+}$. The bound on the VC-dimension follows by construction. ∎
@@
The grid $\mathbb{N}^{n}$ can be viewed as the covering graph of the poset $(\mathbb{N}^{n},\leq)$, where for $x,y\in\mathbb{N}^{n}$ with $x=(x_{1},\ldots,x_{n})$ and $y=(y_{1},\ldots,y_{n})$, we set $x\leq y$ if and only $x_{i}\leq y_{i}$ for $i=1,\ldots,n$. Since ideals of the grid $(\mathbb{N}^{n},\leq)$ are ample, the VC-dimension $\operatorname{dim_{VC}}(L)$ of any such ideal $L$ is the largest subcube of $L$. Notice that $\operatorname{dim_{VC}}(L)$ can be smaller than the dimension $n$ of the grid. It is obvious that there exists a bijection between the ideals of the grid $(\mathbb{N}^{n},\leq)$ and the multisimplicial complexes on a set $X$ of size $n$. The simplicial complexes are in bijection with the ideals of the $n$-cube $\{0,1\}^{n}\subset\mathbb{N}^{n}$. The VC-dimension of any (multi)simplicial complex coincides with its dimension. Hence, we can apply Proposition 22 to obtain the following result:
@@
Every multisimplicial complex of dimension $d$ and $k$ facets is an ideal of an ideal alignment of VC-dimension at most $kd$.
We get an analogous bound for median set systems:
@@
Every median set system of VC-dimension $d$ and having $k$ maximal elements is an ideal of an ideal alignment of VC-dimension at most $kd$.

### v2

@@ 4.3 Convex dimension
Edelman and Jamison [13] provided a nice characterization of convex geometries via order convexity. Given a universe $U=\{e_{1},\ldots,e_{n}\}$, a total order $<$ on $U$, and the reflexive closure $\leq$ of $<$, call an ending interval of $\leq$ any set of the form $\{e\in U:e_{i}\leq e\}$ for some $e_{i}\in U$. Given a set $\Upsilon=\{\leq_{1},\ldots,\leq_{d}\}$ of $d$ total orders on $U$, we say that a set family $\mathcal{C}\subseteq 2^{U}$ is generated by the set $\Upsilon$ if $\varnothing\in\mathcal{C}$ and a nonempty set $C$ belongs to $\mathcal{C}$ if and only if $C$ is the intersection of $d$ ending intervals, one from each order $\leq_{i}$ of $\Upsilon$, i.e., if there exists $(e_{i})_{1\leq i\leq d}\in U^{d}$ such that $C=\{e\in U:e_{i}\leq_{i}e,\forall 1\leq i\leq d\}$. Edelman and Jamison proved the following result:
Theorem 19 ([13, Theorem 5.2]).
@@
An ideal $I$ of a poset is a subset that is downwards-closed, i.e., if $x\leq y$ and $y\in I$, then $x\in I$. From Theorem 20 by definition we get a justification of the name ideals of convex geometries:
@@ 4.5 Examples of ideals of convex geometries
Edelman and Jamison [13] and Korte et al. [24] presented numerous examples of convex geometries and antimatroids, arising from geometry, language theory, and chip firing games. Bandelt et al. [4] presented simplicial complexes and median set systems as examples of bouquets of convex geometries. We show that multisimplicial complexes and median set systems are ideals of particular convex geometries, called downset alignments.
A simplicial complex $L$ on a set $U$ is a family of subsets of $U$, called simplices or faces of $L$, such that if $\sigma\in L$ and $\sigma^{\prime}\subseteq\sigma$, then $\sigma^{\prime}\in L$. The facets of $L$ are the maximal (by inclusion) faces of $L$. The dimension $d$ of $L$ is the size of its largest face minus one. A multi-subset $\sigma$ of $U$ is a subset of $U$ such that each element $e\in U$ is given with its multiplicity $n_{\sigma}(e)$ (the number of times, $e$ occurs in $\sigma$). A multisimplicial complex $L$ on a set $U$ is a family of multi-subsets of $U$, such that if $\sigma\in L$ and $\sigma^{\prime}$ is a multi-subset of $U$ such that $n_{\sigma^{\prime}}(e)\leq n_{\sigma}(e)$ for each $e\in U$, then $\sigma^{\prime}\in L$. The size of a face $\sigma$ of $L$ is the size of $\{e\in U:n_{\sigma}(e)>0\}$. The dimension $d$ of $L$ is the size of its largest face minus one.
A median set system is a set family $(U,\mathcal{C})$ satisfying (C1), (C2), and
@@
The resulting meet-semilattice is called a median semilattice [3, 6, 30]. Median semilattices are locally lower distributive. Indeed, by a result of Birkhoff and Kiss [6] they can be characterized by the property that all principal ideals $\operatorname{\downarrow}x$ are distributive lattices and three elements have an upper bound whenever each pair of them does.
@@
Let $\leq$ be a partial order on $U$. The downset alignment ${\mathcal{D}}_{P}$ [13] of the poset $P=(U,\leq)$ consists of all ideals of $P$. It was noticed in [13] that the downset alignments are convex geometries and it was shown in [13, Theorem 3.2] that a convex geometry $\mathcal{A}=(U,\mathcal{C})$ is a downset alignment if and only if $\mathcal{C}$ is union-closed. From the lattice point of view, downset alignments simply correspond to distributive lattices.
We now define bouquets of downset alignments in the same vein as bouquets of convex geometries were defined. We say that a pair $\mathcal{A}=(U,\mathcal{C})$ is a bouquet of downset alignments if $\mathcal{C}$ satisfies (C1), (C2), (C3), and
  • (C8)
    for all $X,Y\in\mathcal{C}$, if there exists $Z\in\mathcal{C}$ such that $X\cup Y\subseteq Z$, then $X\cup Y\in\mathcal{C}$. (locally union-closed)
Note that bouquets of downset alignments are precisely bouquets of convex geometries with the additional property of being locally union-closed. Note also that principal ideals of bouquets of downset alignments are downset alignments and that ideals of downset alignments are bouquets of downset alignments. Differently from bouquets of convex geometries, that are not always ideals of convex geometries by Proposition 4, the next result shows that bouquets of downset alignments and ideals of downset alignments coincide:
Theorem 23.
Every bouquet of downset alignments $\mathcal{A}=(U,\mathcal{C})$ with $\ell$ maximal convex sets is an ideal of a downset alignment $\mathcal{A}^{+}$ such that $\operatorname{VC-dim}(\mathcal{A}^{+})\leq\ell\cdot\operatorname{VC-dim}(\mathcal{C})$.
@@
Consider a bouquet of downset alignments $\mathcal{A}=(U,\mathcal{C})$. Let $X_{1},\ldots,X_{\ell}$ be the maximal convex sets of $\mathcal{A}$ and let $\operatorname{\downarrow}X_{1},\ldots,\operatorname{\downarrow}X_{\ell}$ be their principal ideals. For each $1\leq i\leq\ell$, let $d_{i}=\operatorname{VC-dim}(\operatorname{\downarrow}X_{i})$. We show by induction on $\ell$ that $\mathcal{A}$ is an ideal of a downset alignment $\mathcal{A}^{+}$ such that $\operatorname{VC-dim}(\mathcal{A}^{+})\leq\sum_{i=1}^{\ell}d_{i}$.
If $\ell=1$, then $\mathcal{A}$ is a downset alignment and we are done. Otherwise, take an element $M\in\mathcal{C}$ that can be written as the intersection of at least two maximal convex sets of $\mathcal{A}$, and that is inclusion maximal for this property. Let $\mathcal{S}$ be the set of all maximal convex sets of $\mathcal{A}$ containing $M$, and assume without loss of generality, that $\mathcal{S}=\{X_{1},\ldots,X_{k}\}$. Consequently, $M=\bigcap_{i=1}^{k}X_{i}$. Let $\mathcal{C}^{\prime}_{0}$ be the set of all sets of the form $Y_{1}\cup\ldots\cup Y_{k}$ where for each $0\leq i\leq k$, $Y_{i}\in\mathcal{C}$ and $M\subseteq Y_{i}\subseteq X_{i}$. Now define $\mathcal{C}^{\prime}$ as $\mathcal{C}\cup\mathcal{C}^{\prime}_{0}$ and set $\mathcal{A}^{\prime}=(U,\mathcal{C}^{\prime})$. Observe that the maximal convex sets of $\mathcal{A}^{\prime}$ are the sets $X_{k+1},\ldots,X_{\ell}$, and $X^{\prime}=\bigcup_{i=1}^{k}X_{i}$.
@@
We first show that $\mathcal{C}^{\prime}$ is still a bouquet of downset alignments, i.e., that $\mathcal{C}^{\prime}$ satisfies (C1), (C2), (C3) and (C8). Observe that $\varnothing\in\mathcal{C}\subseteq\mathcal{C}^{\prime}$, establishing (C1). To prove that $\mathcal{C}^{\prime}$ is intersection-closed, let $A,B\in\mathcal{C}^{\prime}$. If $A,B\in\mathcal{C}$, then there is nothing to show, so suppose first $A\in\mathcal{C}$ and $B\in\mathcal{C}^{\prime}\setminus\mathcal{C}$. Hence we can represent $B=Y_{1}\cup\ldots\cup Y_{k}$ and we consider $A\cap(Y_{1}\cup\ldots\cup Y_{k})=(A\cap Y_{1})\cup\ldots\cup(A\cap Y_{k})$. Since all of these sets are subsets of $A$, since $\mathcal{C}$ is locally union-closed, their union is in $\mathcal{C}$. If $A,B\in\mathcal{C}^{\prime}\setminus\mathcal{C}$, then $A=Y_{1}\cup\ldots\cup Y_{k}$ and $B=Y^{\prime}_{1}\cup\ldots\cup Y^{\prime}_{k}$ and $(Y_{1}\cup\ldots\cup Y_{k})\cap(Y^{\prime}_{1}\cup\ldots\cup Y^{\prime}_{k})=Y_{1}\cap(Y^{\prime}_{1}\cup\ldots\cup Y^{\prime}_{k})\cup\ldots Y_{k}\cap(Y^{\prime}_{1}\cup\ldots\cup Y^{\prime}_{k})$ but for each $i$ here we have $M\subseteq Y_{i}\cap(Y^{\prime}_{1}\cup\ldots\cup Y^{\prime}_{k})\subseteq X_{i}$, hence their union is in $\mathcal{C}^{\prime}$ by definition of $\mathcal{C}^{\prime}$.
@@
In order to show that $\mathcal{C}^{\prime}$ is locally union-closed let $A,B\in\mathcal{C}^{\prime}$ and let $A,B\subseteq Z$. Since we can assume that $A\notin\mathcal{C}$, without loss of generality we can set $Z=\bigcup_{i=1}^{k}X_{i}$ and $A=Y_{1}\cup\ldots\cup Y_{k}$ as before. Now, $B=(B\cap X_{1})\cup\ldots\cup(B\cap X_{k})$, where each term is in $\mathcal{C}$. Then, $A\cup B=(Y_{1}\cup(B\cap X_{1})\cup\ldots\cup Y_{k}\cup(B\cap X_{k}))$, where each term is in $\mathcal{C}$ by local union-closedness and each term contains $M$ as a subset. Hence by definition of $\mathcal{C}^{\prime}$, $A\cup B\in\mathcal{C}^{\prime}$.
@@
Now, let us show that $\mathcal{C}^{\prime}$ is locally extendable. For this purpose let $A,B\in\mathcal{C}^{\prime}$ be distinct sets such that $B$ is a maximal convex set and $A\subseteq B$. If $B\in\mathcal{C}$, then since $\operatorname{\downarrow}B$ does not contains any set of $\mathcal{C}^{\prime}_{0}$, also $A\in\mathcal{C}$ and there is nothing to show. So suppose that $B\in\mathcal{C}^{\prime}\setminus\mathcal{C}$, i.e., that $B=X^{\prime}=\bigcup_{i=1}^{k}X_{i}$ (since $B$ is a maximal convex set of $\mathcal{C}^{\prime}$). If $A\in\mathcal{C}$, then $A\cap X_{i}\subsetneq X_{i}$ for some $1\leq i\leq k$, and thus, there is an element $e\in X_{i}\setminus A\subseteq B\setminus A$ such that $(A\cap X_{i})\cup e\in\mathcal{C}$ since $\mathcal{C}$ is locally extendable. Now, $\mathcal{C}^{\prime}$ is locally union-closed also $A\cup(A\cap X_{i})\cup e=A\cup e\in\mathcal{C}^{\prime}$. If $A\in\mathcal{C}^{\prime}\setminus\mathcal{C}$, then $A=Y_{1}\cup\ldots\cup Y_{k}$ with $Y_{i}\subseteq X_{i}$ for $1\leq i\leq k$ and $Y_{j}\subsetneq X_{j}$ for some $1\leq j\leq k$ (since $A\subsetneq B=X^{\prime}$). Since $\mathcal{C}$ is locally extendable, there is $e\in X_{j}\setminus Y_{j}\subseteq B\setminus A$ such that $Y_{j}\cup e\in\mathcal{C}$. But then by definition of $\mathcal{C}^{\prime}$, we have that $A\cup e\in\mathcal{C}^{\prime}$.
Let us now show that $\mathcal{C}$ is an ideal of $\mathcal{C}^{\prime}$. For this, suppose there exist $X\in\mathcal{C}$ and $R\in\mathcal{C}^{\prime}_{0}=\mathcal{C}^{\prime}\setminus\mathcal{C}$ such that $R\subseteq X$. We can assume that $X$ is a maximal convex set and since $M\subseteq R$ by definition of $\mathcal{C}^{\prime}_{0}$, $X$ necessarily belongs to $\mathcal{S}$ and we can assume that $X=X_{1}$. Let $R=Y_{1}\cup\ldots\cup Y_{k}\subseteq X_{1}$ where for all $1\leq i\leq k$, $M\subseteq Y_{i}\subseteq X_{i}$ and $Y_{i}\in C$. Since $R\notin\mathcal{C}$, necessarily $Y_{1}\subsetneq R$ and thus we can assume that $M\subsetneq Y_{2}$. Since, $Y_{2}\subseteq R\subseteq X_{1}$ and $Y_{2}\subseteq X_{2}$, we have $M\subsetneq Y_{2}\subseteq X_{1}\cap X_{2}$, contradicting the maximality of $M$.
We now show that $\operatorname{VC-dim}(\operatorname{\downarrow}X^{\prime})\leq\sum_{i={1}}^{k}\operatorname{VC-dim}(\operatorname{\downarrow}X_{i})$. Consider a subset $Z\subseteq X^{\prime}$ with $|Z|=\operatorname{VC-dim}(\operatorname{\downarrow}X^{\prime})$ that is shattered by $\operatorname{\downarrow}X^{\prime}$. For each $1\leq i\leq k$, let $Z_{i}=X\cap X_{i}$. We claim that $Z_{i}$ is shattered by $\operatorname{\downarrow}X_{i}$. Indeed, for any $Z^{\prime\prime}\subseteq Z_{i}\subseteq Z$, there exists $X^{\prime\prime}\in\operatorname{\downarrow}X^{\prime}$ such that $X^{\prime\prime}\cap Z=Z^{\prime\prime}$. Observe that $(X^{\prime\prime}\cap X_{i})\cap Z_{i}=X^{\prime\prime}\cap X_{i}\cap Z=Z^{\prime\prime}$ since $X^{\prime\prime}\cap Z=Z^{\prime\prime}\subseteq Z_{i}\subseteq X_{i}$. Note also that $X^{\prime\prime}\cap X_{i}\in\mathcal{C}^{\prime}$ and thus $X^{\prime\prime}\cap X_{i}\in\operatorname{\downarrow}X_{i}$. Consequently, $Z_{i}$ is shattered by $\operatorname{\downarrow}X_{i}$. Therefore, since $Z=\cup_{i=1}^{k}Z_{i}$, $\operatorname{VC-dim}(\operatorname{\downarrow}X^{\prime})=|Z|\leq\sum_{i=1}^{k}|Z_{i}|\leq\sum_{i=1}^{k}\operatorname{VC-dim}(\operatorname{\downarrow}X_{i})$, and we are done.
Consequently, $\mathcal{A}^{\prime}$ is a bouquet of downset alignments, $\mathcal{A}$ is an ideal of $\mathcal{A}^{\prime}$, and the maximal convex sets of $\mathcal{A}^{\prime}$ are the sets $X_{k+1},\ldots,X_{\ell}$, and $X^{\prime}=\bigcup_{i=1}^{k}X_{i}$. Since $k\geq 2$, $\mathcal{C}^{\prime}$ has less inclusion-maximal convex sets than $\mathcal{C}$. By induction hypothesis, $\mathcal{A}^{\prime}$ is an ideal of a downset alignment $\mathcal{A}^{+}$ such that $\operatorname{VC-dim}(\mathcal{A}^{+})\leq\operatorname{VC-dim}(\operatorname{\downarrow}X^{\prime})+\sum_{i={k+1}}^{\ell}\operatorname{VC-dim}(\operatorname{\downarrow}X_{i})\leq\sum_{i={1}}^{k}\operatorname{VC-dim}(\operatorname{\downarrow}X_{i})+\sum_{i={k+1}}^{\ell}\operatorname{VC-dim}(\operatorname{\downarrow}X_{i})=\sum_{i={1}}^{\ell}\operatorname{VC-dim}(\operatorname{\downarrow}X_{i})$. By transitivity, $\mathcal{A}$ is an ideal of $\mathcal{A}^{+}$. This concludes the proof. ∎
@@
The grid $\mathbb{N}^{n}$ can be viewed as the covering graph of the poset $(\mathbb{N}^{n},\leq)$, where for $x,y\in\mathbb{N}^{n}$ with $x=(x_{1},\ldots,x_{n})$ and $y=(y_{1},\ldots,y_{n})$, we set $x\leq y$ if and only $x_{i}\leq y_{i}$ for $i=1,\ldots,n$. Since ideals of the grid $(\mathbb{N}^{n},\leq)$ are ample, the VC-dimension $\operatorname{dim_{VC}}(L)$ of any such ideal $L$ is the largest subcube of $L$. Notice that $\operatorname{dim_{VC}}(L)$ can be smaller than the dimension $n$ of the grid. It is obvious that there exists a bijection between the ideals of the grid $(\mathbb{N}^{n},\leq)$ and the multisimplicial complexes on a set $X$ of size $n$. The simplicial complexes are in bijection with the ideals of the $n$-cube $\{0,1\}^{n}\subset\mathbb{N}^{n}$. The VC-dimension of any (multi)simplicial complex coincides with its dimension. Hence, we can apply Theorem 23 to obtain the following result:
@@
Every multisimplicial complex of dimension $d$ and $\ell$ facets is an ideal of a downset alignment of VC-dimension at most $\ell d$.
We get an analogous bound for median set systems (since they principal ideals are distributive lattices, median set systems are bouquets of downset alignments):
@@
Every median set system of VC-dimension $d$ and having $\ell$ maximal elements is an ideal of a downset alignment of VC-dimension at most $\ell d$.